OpenStudy (anonymous):

Can someone show me how to work this out? 32x^3-28x^2-87x-27 divide by 8x+3

OpenStudy (anonymous):

Are you studying long division?

OpenStudy (anonymous):

Do you know synthetic division?

OpenStudy (anonymous):

yes, long division, factoring polynomials

OpenStudy (anonymous):

Just a minute

OpenStudy (anonymous):

How many times does 8x go into 32x^3 or you can ask yourself what is 32x^3 divided by 8x?

OpenStudy (anonymous):

would it be 4x^3, but how do if figure out the rest -28x^2-87x-27?

OpenStudy (anonymous):

It would be 4x^2 because you already have 1 of the x's with the 8x 8x(4x^2) = 32x^3 This is a long process... so... put 4x^2 above the 32x^3 and multiply 4x^2 (8x+3) and put that under the polynomial

OpenStudy (anonymous):

4x^2 ------------------------ 8x + 3 | 32x^3 - 28x^2 - 87x - 27 32x^3 + 12x^2

OpenStudy (anonymous):

Now draw a line under the bottom part, change the signs 32x^3 + 12x^2 becomes -32x^3 - 12x^2 and then add. you should end up with -40x^2

OpenStudy (anonymous):

oh ok, it is easier to "see" this way..thank you

OpenStudy (anonymous):

4x^2 ------------------------ 8x + 3 | 32x^3 - 28x^2 - 87x - 27 -32x^3 - 12x^2 -------------- -40x^2 = 87x then you ask yourself 8x times what is -40x^2 and put that above the -28x^2 and repeat the process

OpenStudy (radar):

India, have you figured out what to multiply 8x to get -40x^2? Don't give up.

OpenStudy (anonymous):

Sorry that is a - 87 not an = 87x

OpenStudy (radar):

That polynomial will factor to 3 binomials.

OpenStudy (radar):

Earth to India, have you quit?

OpenStudy (radar):

(4x-9)(x+1)(8x+3)

OpenStudy (radar):

You need to practice the long division method.

OpenStudy (anonymous):

Thank you, I understood blexting explained it. Thanks again

OpenStudy (radar):

Oh O.K. I thought you might of got stuck on the 8x times ? =40x^2. Good to hear you crossed that hurdle. Good luck